1)     When price of X commodity rises from Br. 10 to 15, demand for

Y commodity declines from 200 to 150.

A) Calculate cross price elasticity.

B) Based on the result, what kind of relation exists between the two goods?


1
Expert's answer
2022-01-16T13:31:22-0500

Solution:

A.). Cross-price elasticity of demand = "=\frac{\%\;change\; in\; quantity\; demanded \;of\; Y}{\%\; change\; in\; the \;price \;for\; X }"


="\frac{150 -200}{(150+200)/2 } \div \frac{15 -10}{(15+10)/2 } = \frac{-50}{175} \div\frac{5}{12.5} = \frac{-0.2857}{0.4} = -0.71"


Cross-price elasticity of demand = -0.71

 

B.). Since the cross-price elasticity of demand is negative, we can conclude that good X and Y are complementary goods. That is when the price of commodity X rises, consumers will purchase less of commodity X, thus purchasing less commodity Y and vice versa.